Upendo ltd wishes to take advantage of the new commercial paper now popular. It


wishes to issue two debenture papers. Both bear coupons of 12% and the effective yield


required on each is 20%. Paper A has a maturity of 10 years and paper B a maturity of


20 years. Both will be paying interest annually and Kshs. 100,000 at maturity. What is


the price of each paper? If the effective yield on each paper rises to 24%, what is the


price of each paper?

1
Expert's answer
2021-11-02T18:59:03-0400
Cost of Debentures=Interest(1-t) + (Par Value - Market Value)/0.5(Par Value-Market Value)
         Ke={D0(1+g)/(Market Price per share)}+g
                ={12(1+24)/20}+24
                     15+24=39

Cost of debentures(A)={12(1-0)+(24-12)1/10}/{1/2(24+12)}

                        =(12+1.82)/18
                               =0.7678
                               =76.78%

Cost of debenture(B)={12(1-0)+(24-12)1/20}/{0.5(24+12)}

                         =1.132/18
                           =0.0628
                             =62.8%